How to find the sum of sigma with the formula being 3an -n, and the starting and ending points being one and five, respectively?
I approached the problem by setting the formula equal to zero and solving for an, which gives me an = -n/3, ultimately getting a sum of -5.
Is this right at all or am I way off?
Thank you so much.

Answer by stanbon(75887)

How to find the sum of sigma with the formula being 3an -n, and the starting and ending points being one and five, respectively?
I approached the problem by setting the formula equal to zero and solving for an, which gives me an = -n/3, ultimately getting a sum of -5.
-----------------------
If a(n) = -n/3
n=1: a(1) = -1/3
n=2: a(2) = -2/3
n=3: a(3) = -3/3
n=4: a(4) = -4/3
n=5: a(5) = -5/3
------------------------
Adding you would get: [-1-2-3-4-5]/3 = -15/3 = -5
=====
Your answer looks good.
Cheers,
Stan H.